39 packages are randomly selected from packages received by a parcel service. The sample has a mean weight of 15.3 pounds and a standard deviation of 1.9 pounds. What is the 95 percent confidence interval for the true mean weight, µ, of all packages received by the parcel service?
Answer by Boreal(15235)

The 95% half-interval is t0.975, df=38*1.9/sqrt(39)
t=2.02
half-interval is 2.02*1.9/sqrt(39)=0.61
(14.69, 15.91)